if the numerator of a certain fraction is increased by 6 and its denominator is decreased by 5, the resulting fraction is equal to 3/4. if the reciprocal of the original fraction is decreased by 1, the resulting fraction is 16/9. find the original fraction.: if the numerator of a certain fraction is increased by 6 and its denominator is decreased by 5, the resulting fraction is equal to 3/4. if the reciprocal of the original fraction is decreased by 1, the resulting fraction is 16/9. find the original fraction.
Answer by ankor@dixie-net.com(4489)

if the numerator of a certain fraction is increased by 6 and its denominator is decreased by 5, the resulting fraction is equal to 3/4. if the reciprocal of the original fraction is decreased by 1, the resulting fraction is 16/9. find the original fraction.
:
Let "a certain fraction" = n/d
and
"if the numerator of a certain fraction is increased by 6 and its denominator is decreased by 5, the resulting fraction is equal to 3/4."
((n+6))/((d-5)) = 3/4
and
"If the reciprocal of the original fraction is decreased by 1, the resulting fraction is 16/9."
d/n - 1 = 16/9
Add 1 to both sides:
d/n = 16/9 + 9/9
d/n = 25/9
That means:
n = 9
-- --- is the original fraction
d = 25
:
Using the 1st statment to check this:
((9+6))/((25-5)) = 15/20 = 3/4